# Towards Principled Superhuman AI for Multiplayer Symmetric Games

### v1

@@ Abstract
Multiplayer games, when the number of players exceeds two, present unique challenges that fundamentally distinguish them from the extensively studied two-player zero-sum games. These challenges arise from the non-uniqueness of equilibria and the risk of agents performing highly suboptimally when adopting equilibrium strategies. While a line of recent works developed learning systems successfully achieving human-level or even superhuman performance in popular multiplayer games such as Mahjong, Poker, and Diplomacy, two critical questions remain unaddressed: (1) What is the correct solution concept that AI agents should find? and (2) What is the general algorithmic framework that provably solves all games within this class? This paper takes the first step towards solving these unique challenges of multiplayer games by provably addressing both questions in multiplayer symmetric normal-form games. We also demonstrate that many meta-algorithms developed in prior practical systems for multiplayer games can fail to achieve even the basic goal of obtaining agent’s equal share of the total reward.
@@ 1 Introduction
In recent years, AI has achieved remarkable success in multi-agent decision-making problems, particularly in a wide range of strategic games. These include, but are not limited to, Go (Silver et al.,, 2016), Mahjong (Li et al.,, 2020), Poker (Moravčík et al.,, 2017; Brown and Sandholm,, 2018, 2019), Starcraft 2 (Vinyals et al.,, 2019), DOTA 2 (Berner et al.,, 2019), League of Legends (Ye et al.,, 2020), and Diplomacy (Gray et al.,, 2020; Bakhtin et al.,, 2022; † et al., 2022, FAIR). The majority of these games are two-player zero-sum games [^1] Games such as DOTA and League of Legends, despite involving two teams, can be mostly considered similar to two-player zero-sum games in terms of their game structures and solutions., where Nash equilibria always exist and can be computed efficiently. Nash equilibria in two-player zero-sum games are also non-exploitable—an agent employing a Nash equilibrium policy will not lose even when facing an adversarial opponent who seeks to exploit the agent’s weaknesses. Although such equilibrium strategies do not necessarily capitalize on opponents’ weaknesses or guarantee large-margin victories, human players often adopt suboptimal policies that deviate significantly from equilibria in complex games with large state spaces. Consequently, AI agents who adopt equilibrium strategies often outperform humans in practice for two-player zero-sum games.
@@
In contrast, multiplayer games—defined here as those with more than two players—exhibit fundamentally different game structures compared to two-player zero-sum games. This distinction introduces several unique challenges. Firstly, Nash equilibria are no longer efficiently computable (Daskalakis et al.,, 2009; Chen and Deng,, 2005). Moreover, there may exist multiple Nash equilibria with distinct values. Such non-uniqueness in equilibria raises a critical concern about the adoption of equilibrium strategies in multiplayer settings: if an AI agent adopts an equilibrium that is different from other players, collectively, they are not playing any single equilibrium, which undermines the equilibrium property that dissuades agents from changing strategies as long as others maintain theirs. Finally, in multiplayer games, equilibrium strategies are no longer non-exploitable. Although the introduction of alternative equilibrium notions such as Correlated Equilibria (CE) or Coarse Correlated Equilibria (CCE) alleviates computational hardness, issues of non-uniqueness and potential exploitability remain. This leads to the first critical question:
What is the correct solution concept for an AI agent to learn in multiplayer games?
@@
While not answering this question directly, several practical systems for multiplayer games have already achieved human-level or even superhuman performance in several popular games including Mahjong (Li et al.,, 2020) (4 players), Poker (Brown and Sandholm,, 2019) (6 players), and Diplomacy (Bakhtin et al.,, 2022; † et al., 2022, FAIR) (7 players). These systems focus on developing algorithmic frameworks capable of learning effective strategies that excel in competitive settings, such as ladders, online gaming platforms, or tournaments against human players. Generally, most of these systems rely on a basic self-play framework, starting from scratch or from human policies acquired through behavior cloning, with or without regularizations. Since all these systems demand substantial computational power, significant human expertise, and extensive engineering efforts, the general applicability of the algorithmic frameworks developed in these studies remains to be tested in a wide range of multiplayer games beyond Mahjong, Poker, or Diplomacy. This leads to the second critical question:
What is the general algorithmic framework that can provably solve all multiplayer games within certain rich classes?
@@
In this paper, we consider an important subclass of multiplayer games: symmetric constant-sum games, which are prevalent in games involving more than two players. Examples include previously discussed multiplayer games like Mahjong, Poker, Diplomacy, as well as a variety of board games such as Avalon (Light et al.,, 2023), Mafia and Catan. [^2] All these examples are symmetric games up to randomization of the seating. Symmetry brings fairness among players, and a natural baseline for symmetric constant-sum games is that the AI agent should, at a minimum, secure an equal share of the total reward. For instance, in a four-player game with only one winner, the AI should aim to win at least one-fourth of the games. This paper also focuses on the formulation of normal-form games, which are not only foundational but also expressive enough to encapsulate sequential games like extensive-form games or Markov games as special cases. [^3] Extensive-Form Games (EFGs) or Markov Games (MGs) can be viewed as special cases of normal-form games, where each action in normal-form games corresponds to a policy in EFGs or MGs, although such representations may not always be efficient.
This paper takes an initial step towards addressing the grand challenges of general multiplayer games by effectively tackling the two highlighted questions within the context of multiplayer symmetric normal-form games. Regarding the first question of solution concepts, we discuss the inadequacy of classical equilibrium solutions in meeting the baseline requirement of achieving an equal share in symmetric games. We further establish important claims that (1) as long as opponents can deploy different individual policies, even playing their best response can fail to achieve an equal share; (2) there may not exist a non-exploitable strategy even when all opponents are restricted to play the same strategy (See Section 4 for details). This unfortunately leaves us with the only viable option that the agent must learn strategies that adapt to the identical strategy played by the opponents. Solution concepts that fall into this category, such as the best response to the identical strategy of the opponents, always meet the baseline requirement. While assuming all opponents playing identical strategy appears restrictive, we argue that it has already been implicitly adopted in many modern AI systems for multiplayer games. We also prove that such an assumption approximately holds in multiplayer gaming platforms with a large player base.
For the second question concerning general algorithms, this paper illustrates how we can leverage existing no-regret and no-dynamic regret learning tools appropriately to meet the baseline requirement of equal share with regret guarantees, particularly when human policies are either stationary or evolving slowly. We also present matching lower bounds that show these regret guarantees cannot be significantly improved in the worst case. Our experimental results demonstrate that while our algorithmic solutions consistently outperform static human policies, the self-play meta-algorithms developed in previous state-of-the-art systems for multiplayer games like Mahjong, Poker, and Diplomacy may converge to suboptimal solutions, resulting in the AI agent consistently losing the game even against non-adaptive human players.
@@ 1.1 Related work
Human-level or superhuman AI in practice
Building superhuman AI has long been a goal in various games. A large body of works in this line focus on two-player or two-team zero-sum games like Chess (Campbell et al.,, 2002), Go (Silver et al.,, 2016), Heads-Up Texas Hold’em (Moravčík et al.,, 2017; Brown and Sandholm,, 2018), Starcraft 2 (Vinyals et al.,, 2019), DOTA 2 (Berner et al.,, 2019) and League of Legends (Ye et al.,, 2020). Most of them are based on finding equilibria via self-play, fictitious play, league training, etc. There is comparatively much less amount of work on games with more than two players, whose game structures are fundamentally different from two-player zero-sum games. Several remarkable multiplayer successes include Poker (Brown and Sandholm,, 2019), Mahjong (Li et al.,, 2020), Doudizhu (Zha et al.,, 2021) and Diplomacy (Bakhtin et al.,, 2022; † et al., 2022, FAIR). Despite lacking a clearly formulated learning objective, these works typically design meta-algorithms, which include initially training the model using behavior cloning from human players, then enhancing it through self-play, and finally applying adaptations based on the game’s specific structure or human expertise. It remains elusive whether such a recipe is generally effective for a wide range of multiplayer games.
Existing results on symmetric games
Von Neumann & Morgenstern, 1947 (Von Neumann and Morgenstern,, 1947) gave the first definition of symmetric games and used the three-player majority-vote example to showcase the stark difference between symmetric three-player zero-sum games and symmetric two-player zero-sum games. In his seminal paper that introduced Nash equilibrium, Nash proved that a symmetric finite multi-player game must have a symmetric Nash equilibrium (Nash,, 1951). However, this existence result might mean little from an individual standpoint, as there is no reason a priori to assume that other players are indeed playing according to this symmetric equilibrium. Papadimitriou & Roughgarden, 2005(Papadimitriou and Roughgarden,, 2005) studied the computational complexity of finding the Nash equilibrium in symmetric multi-player games when the number of actions available is much smaller than the number of players, and gave a polynomial-time algorithm for the problem. In this case, symmetry greatly reduced the computational complexity (as computing Nash in general is PPAD-hard). Daskalakis et al., 2009 (Daskalakis,, 2009) proposed anonymous games, a generalization of symmetric games.
@@
There is a rich literature on applying no-regret learning algorithms to learning equilibria in games. It is well-known that if all agents have no-regret, the resulting empirical average would be an approximate Coarse Correlated Equilibrium (CCE) (Young,, 2004), while if all agents have no swap-regret, the resulting empirical average would be an $\epsilon$-Correlated Equilibrium (CE) (Hart and Mas-Colell,, 2000; Cesa-Bianchi and Lugosi,, 2006). Later work continuing this line of research includes those with faster convergence rates (Syrgkanis et al.,, 2015; Chen and Peng,, 2020; Daskalakis et al.,, 2021), last-iterate convergence guarantees (Daskalakis and Panageas,, 2018; Wei et al.,, 2020), and extension to extensive-form games (Celli et al.,, 2020; Bai et al., 2022b, ; Bai et al., 2022a, ; Song et al.,, 2022) and Markov games (Song et al.,, 2021; Jin et al.,, 2021).
@@ 2 Preliminaries
An $n$-player normal-form game consists of a finite set of $n$ players, where each player has an action space $\mathcal{A}_{i}$ and a corresponding payoff function $U_{i}:\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to[-1,1]$ with $U_{i}(a_{1},\ldots,a_{n})$ denotes the payoff received by the $i$-th player if $n$ players are taking joint actions $(a_{1},\ldots,a_{n})$. This paper focuses on symmetric and zero-sum normal-form games. [^4] Our work trivially extends to symmetric and constant-sum normal-form games, since one can convert constant-sum games to zero-sum games by offsetting all players’ payoff functions by a constant.
Definition 2.1 (Symmetric zero-sum normal-form game).
For an $n$-player normal-form game with an action space $\mathcal{A}_{i}$ and a payoff $U_{i}$ for player $i$, we say the game is symmetric if (1) $\mathcal{A}_{i}=\mathcal{A}$, for all $i\in[n]$; (2) for any permutation $\pi$: $U_{i}(a_{1},\cdots,a_{n})=U_{\pi^{-1}(i)}(a_{\pi(1)},\cdots,a_{\pi(n)})$. We say the game is zero-sum if $\sum^{n}_{i=1}U_{i}(a)=0$ for all $a\in\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$.
Briefly speaking, in a symmetric game, the payoffs for employing a specific action are determined solely by the actions used by others, agnostic of the identities of the players using them. Thus, the payoff function of the first player, denoted as $U_{1}$, is sufficient to encapsulate the entire game. For $i\in[n]$, $x_{i}\in\Delta(\mathcal{A})$ denotes a mixed strategy of the $i$-th player and $x_{-i}\in\Delta(\mathcal{A}^{\otimes n-1})$ denotes a mixed strategy of the other players. Correspondingly, $a_{i}\in\mathcal{A}$ denotes the action of the $i$-th player and $a_{-i}\in\mathcal{A}^{\otimes n-1}$ denotes the action of the other players. For any $a_{i}\in\mathcal{A}$, we denote $U_{i}(a_{i},x_{-i}):=\mathbb{E}_{a_{-i}\sim x_{-i}}\left[U_{i}(a_{i},a_{-i})\right].$
@@
Nash equilibrium is the most commonly-used solution concept for games: a mixed strategy $x\in\Delta(\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n})$ of all players is said to be Nash Equilibrium (NE) if $x$ is a product distribution, and no player could gain by deviating her own strategy while holding all other players’ strategies fixed. That is, for all $i\in[n]$ and $a^{\prime}_{i}\in\mathcal{A}_{i}$, $\mathbb{E}_{a\sim x}[U_{i}(a_{i},a_{-i})]\geq\mathbb{E}_{a\sim x}[U_{i}(a^{\prime}_{i},a_{-i})]$.
@@
There are also two equilibrium notions relaxing the notion of Nash equilibrium by no longer requiring $x$ to be a product distribution. It allows general joint distribution $x$ which describes correlated policies among players. In particular, (1) $x$ is a Correlated Equilibrium (CE) if for all $i\in[n]$ and $a^{\prime}_{i}\in\mathcal{A}_{i}$, $\mathbb{E}_{a\sim x}[U_{i}(a_{i},a_{-i})\mid a_{i}]\geq\mathbb{E}_{a\sim x}[U_{i}(a^{\prime}_{i},a_{-i})\mid a_{i}]$, and (2) $x$ is Coarse Correlated Equilibrium (CCE) if for all $i\in[n]$ and $a^{\prime}_{i}\in\mathcal{A}_{i}$: $\mathbb{E}_{a\sim x}[U_{i}(a_{i},a_{-i})]\geq\mathbb{E}_{a\sim x}[U_{i}(a^{\prime}_{i},a_{-i})]$. The major difference between those two notions is in the cases when the agent deviates from her current strategy, whether she is still allowed to observe the randomness in drawing actions from the correlated policy. The relationship among various equilibrium concepts is encapsulated by ${\rm NE}\subset{\rm CE}\subset{\rm CCE}$.
@@
We note in two-player zero-sum games, a Nash equilibrium is non-exploitable. In math, if $(\mu^{\star},\nu^{\star})$ is the Nash equilibrium, we have $\min_{\nu}U_{1}(\mu^{\star},\nu)=\max_{\mu}\min_{\nu}U_{1}(\mu,\nu)=0$.
No-regret learning
No-regret learning is a commonly adopted strategy in game theory to find equilibrium solutions. We consider a $T$-step learning procedure, where for each round $t\in[T]$: (1) the agent picks a mixed strategy $\mu^{t}$ over $\mathcal{A}$, (2) the environment picks an adversarial loss $\ell_{t}\in[0,1]^{|\mathcal{A}|}$. The expected utility for $t$-th round is defined as $-\langle\mu^{t},\ell_{t}\rangle$. To measure the performance of a particular algorithm, a common approach is to consider the regret, where the algorithm’s performance is compared against the single best action in hindsight. Specifically, for policy sequence $(\mu^{1},\ldots,\mu^{T})$ taken by an algorithm, we define
${\text{Reg}(T)}:=\sum_{t=1}^{T}\langle\mu^{t},\ell_{t}\rangle-\min_{a\in\mathcal{A}}\sum^{T}_{t=1}\ell_{t}(a),$
$\mu^{t+1}(a)\propto\mu^{t}(a)e^{-\eta_{t}\ell_{t}(a)},\quad\text{for}\quad\forall a\in\mathcal{A}.$ (1)
where $\eta_{t}$ is the learning rate.
Relation between no-regret learning and equilibria
It is known that in the normal-form game setup, if each player independently plays a no-regret learning algorithm treating all other players as adversarial environments, and we denote $x^{t}=x_{1}^{t}\times\cdots\times x_{n}^{t}$ as the joint strategy by all players at round $t$, then we have the average policy $(1/T)\sum_{t=1}^{T}x^{t}$ converges to CCEs when $T$ is large.
3 Equilibria and Self-play are Insufficient for Multiplayer Games
In a multiplayer symmetric zero-sum game, every player is created equal, and the equal share for each player is zero. Intuitively, one might expect a good player to receive a non-negative expected payoff, making this a minimal baseline for a well-designed AI agent. However, in this section, we demonstrate that both the existing equilibria notions and the algorithmic framework of self-play from scratch—widely used in achieving superhuman performance in practical games—are insufficient to secure this seemingly straightforward goal.
To illustrate this, we consider the following $3$-player majority vote game involving an AI agent and two other players.

### v2

# Securing Equal Share: A Principled Approach for Learning Multiplayer Symmetric Games

@@ Abstract
This paper examines multiplayer symmetric constant-sum games with more than two players in a competitive setting, including examples like Mahjong, Poker, and various board and video games. In contrast to two-player zero-sum games, equilibria in multiplayer games are neither unique nor non-exploitable, failing to provide meaningful guarantees when competing against opponents who play different equilibria or non-equilibrium strategies. This gives rise to a series of long-lasting fundamental questions in multiplayer games regarding suitable objectives, solution concepts, and principled algorithms. This paper takes an initial step towards addressing these challenges by focusing on the natural objective of equal share—securing an expected payoff of $C/n$ in an $n$-player symmetric game with a total payoff of $C$. We rigorously identify the theoretical conditions under which achieving an equal share is tractable and design a series of efficient algorithms, inspired by no-regret learning, that provably attain approximate equal share across various settings. Furthermore, we provide complementary lower bounds that justify the sharpness of our theoretical results. Our experimental results highlight worst-case scenarios where meta-algorithms from prior state-of-the-art systems for multiplayer games fail to secure an equal share, while our algorithm succeeds, demonstrating the effectiveness of our approach.
@@ 1 Introduction
In recent years, AI systems have achieved remarkable success in multi-agent decision-making problems, particularly in a wide range of strategic games. These include, but are not limited to, Go (Silver et al., 2016), Mahjong (Li et al., 2020), Poker (Moravčík et al., 2017; Brown & Sandholm, 2018, 2019), Starcraft 2 (Vinyals et al., 2019), DOTA 2 (Berner et al., 2019), League of Legends (Ye et al., 2020), and Diplomacy (Gray et al., 2020; Bakhtin et al., 2022; † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR). Many of these games are two-player zero-sum games [^1] Games such as DOTA and League of Legends, despite involving two teams, can be mostly considered similar to two-player zero-sum games in terms of their game structures and solutions., where Nash equilibria always exist and can be computed in polynomial time. Nash equilibria in two-player zero-sum games are also non-exploitable—an agent employing a Nash equilibrium strategy will not lose even when facing an adversarial opponent who seeks to exploit the agent’s weaknesses. Although such equilibrium strategies do not necessarily capitalize on opponents’ weaknesses or guarantee large-margin victories, human players often adopt suboptimal strategies that deviate significantly from equilibria in complex games with large state spaces. Consequently, AI agents who adopt equilibrium strategies often outperform humans in practice for two-player zero-sum games.
@@
In contrast, multiplayer games—defined here as those with more than two players—exhibit fundamentally different game structures compared to two-player zero-sum games. This distinction introduces several unique challenges. Firstly, Nash equilibria are believed to be no longer computable in polynomial time (Daskalakis et al., 2009; Chen & Deng, 2005). Moreover, there may exist multiple Nash equilibria with distinct values. Such non-uniqueness in equilibria raises a critical concern about the adoption of equilibrium strategies in multiplayer settings: if a learning agent adopts an equilibrium that is different from other players, collectively, they are not playing any single equilibrium, which undermines the equilibrium property that dissuades the agent from changing its strategy as long as others maintain theirs. Finally, in multiplayer games, equilibrium strategies are no longer non-exploitable and fail to provide meaningful guarantees when competing against opponents who are not playing equilibria. Although the introduction of alternative equilibrium notions such as (coarse) correlated equilibria alleviates computational hardness, issues of non-uniqueness and the lack of guarantees in the general settings remain. This leads to the first critical question:
What is the suitable solution concept for learning in multiplayer games?
@@
Due to the presence of such fundamental challenge, even state-of-the-art expert-level or superhuman AI systems for popular multiplayer games, including Diplomacy (Bakhtin et al., 2022; † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR) (7 players), Poker (Brown & Sandholm, 2019) (6 players), and Mahjong (Li et al., 2020) (4 players), are designed with limited theoretical supports. These works focus on developing algorithmic frameworks capable of learning effective strategies that excel in ladders, online gaming platforms, or tournaments against opponents. Generally, most of these systems rely on a basic self-play framework, starting from scratch or from opponents’ strategies acquired through behavior cloning, with or without regularization. While the success of these self-play algorithms is remarkable, their performance has been demonstrated primarily within the specific applications they were designed for, often coupled with human expertise and extensive engineering efforts. It remains unclear whether these algorithms are general-purpose solutions that can be readily applied to multiplayer games beyond Mahjong, Poker, or Diplomacy. This leads to the second important question:
What is the principled algorithm that provable learns a rich class of multiplayer games?
@@
In this paper, we consider multiplayer symmetric constant-sum games, which are prevalent in games involving more than two players. Examples include previously discussed multiplayer games like Mahjong, Poker, and Diplomacy, as well as a variety of board games such as Avalon (Light et al., 2023), Mafia, and Catan. [^2] All these examples are symmetric games up to randomization of the seating. Symmetry brings fairness among players, providing a natural baseline where a learning agent should at least secure an equal share — achieving an expected payoff of $C/n$ in an $n$-player symmetric game with a total payoff of $C$. This paper takes an initial step toward addressing the two fundamental questions highlighted above by focusing on securing an equal share in multiplayer symmetric normal-form games [^3] Extensive-Form Games (EFGs) or Markov Games (MGs) can be viewed as special cases of normal-form games, where each action in normal-form games corresponds to a strategy in EFGs or MGs, although such representations may not always be efficient.. The main contributions are summarized as follows:
1. Regarding the question of solution concepts, we first demonstrate the insufficiency of classical equilibrium concepts and general self-play frameworks (learning from scratch) in achieving an equal share in symmetric games. We then proceed to identify the structural conditions where equal share is achievable. In contrast to two-player zero-sum games, we prove that in order to achieve an equal share in multiplayer games: (1) all opponents need to deploy the same strategy; (2) all opponents must have limited adaptivity, and the learning agent has to model the opponents (See Section 4). We show that without either condition, an equal share can not be attained in the worst case. We prove that our identified conditions apply to practical multiplayer gaming platforms with a large player base. They are also tightly connected to the design of many prior modern multiplayer AI agents.
2. Regarding the question of principled algorithms, this paper illustrates how we can leverage existing tools from no-regret learning and no-dynamic-regret learning communities to achieve equal share with provable guarantees. Concretely, this paper considers several opponent settings: fixed, slowly adapting, and opponents that adapt at intermediate rates. For all cases, we design algorithms that approximately achieve equal share, with an error tolerance of $1/\text{poly}(T)$, where $T$ is the total number of games played. Additionally, we provide matching lower bounds, demonstrating that these guarantees cannot be significantly improved in the worst-case scenario.
3. We further complement our theory by experiments on two basic multiplayer symmetric games. Our experimental results illustrate that (1) the self-play meta-algorithms from prior state-of-the-art systems for multiplayer games can fail to secure an equal share even under favorable settings, while our principled algorithm always succeeds; (2) prior meta-algorithm has no clear advantage over our algorithm on exploitability in the worst case. This indicates that prior self-play algorithms are not general-purpose and highlights the effectiveness of our theoretical framework.
@@ 1.1 Related work
AI gaming agents in practice
Building superhuman AI has long been a goal in various games. A large body of works in this line focus on two-player or two-team zero-sum games like Chess (Campbell et al., 2002), Go (Silver et al., 2016), Heads-Up Texas Hold’em (Moravčík et al., 2017; Brown & Sandholm, 2018), Starcraft 2 (Vinyals et al., 2019), DOTA 2 (Berner et al., 2019) and League of Legends (Ye et al., 2020). Most of them are based on finding equilibria via self-play, fictitious play, league training, etc. There is comparatively much less amount of work on games with more than two players, whose game structures are fundamentally different from two-player zero-sum games. Several remarkable multiplayer successes include Poker (Brown & Sandholm, 2019), Mahjong (Li et al., 2020), Doudizhu (Zha et al., 2021) and Diplomacy (Bakhtin et al., 2022; † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR). Despite lacking a clearly formulated learning objective, these works typically design meta-algorithms, which include initially training the model using behavior cloning from opponents, then enhancing it through self-play, and finally applying adaptations based on the game’s specific structure or human expertise. It remains elusive whether such a recipe is generally effective for a wide range of multiplayer games.
Existing results for symmetric games
Von Neumann & Morgenstern (1947) gave the first definition of symmetric games and used the three-player majority-vote example to showcase the stark difference between symmetric three-player zero-sum games and symmetric two-player zero-sum games. In his seminal paper that introduced Nash equilibrium, Nash proved that a symmetric finite multi-player game must have a symmetric Nash equilibrium  (Nash, 1951). However, this existence result holds little significance from an individual standpoint, as there is no reason a priori to assume that other players are indeed playing according to this symmetric equilibrium. Papadimitriou & Roughgarden (2005) studied the computational complexity of finding the Nash equilibrium in symmetric multi-player games when the number of actions available is much smaller than the number of players and gave a polynomial-time algorithm for the problem. In this case, symmetry greatly reduced the computational complexity (as computing Nash in general is PPAD-hard). Daskalakis (2009) proposed anonymous games, a generalization of symmetric games.
@@
There is a rich literature on applying no-regret learning algorithms to learning equilibria in games. It is well-known that if all agents have no regret, the resulting empirical average would be an approximate Coarse Correlated Equilibrium (CCE) (Young, 2004), while if all agents have no swap-regret, the resulting empirical average would be an $\epsilon$-Correlated Equilibrium (CE) (Hart & Mas-Colell, 2000; Cesa-Bianchi & Lugosi, 2006). Later work continuing this line of research includes those with faster convergence rates (Syrgkanis et al., 2015; Chen & Peng, 2020; Daskalakis et al., 2021), last-iterate convergence guarantees (Daskalakis & Panageas, 2018; Wei et al., 2020), and extension to extensive-form games (Celli et al., 2020; Bai et al., 2022b, a; Song et al., 2022) and Markov games (Song et al., 2021; Jin et al., 2021).
@@ 2.1 Normal-form games and equilibrium
An $n$-player normal-form game consists of a finite set of $n$ players, where each player has an action space $\mathcal{A}_{i}$ and a corresponding payoff function $U_{i}:\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to[-1,1]$ with $U_{i}(a_{1},\ldots,a_{n})$ denotes the payoff received by the $i$-th player if $n$ players are taking joint actions $(a_{1},\ldots,a_{n})$. We define a game as constant-sum if there exists constant $C$ such that $\sum_{i=1}^{n}U_{i}(a_{1},\ldots,a_{n})=C$ for all joint actions. We further denote a game as zero-sum if it is a constant-sum game with a total payoff of $C=0$. Normal-form games can represent a wide range of games as their special cases, including sequential games such as extensive-form or Markov games.
Strategy
A (mixed) strategy of a player is a probability distribution over the player’s actions. For $i\in[n]$, we use $a_{i}\in\mathcal{A}$ and $x_{i}\in\Delta(\mathcal{A})$ to denote an action and a mixed strategy of the $i$-th player respectively. We use $a_{-i}\in\mathcal{A}^{\otimes n-1}$ and $x_{-i}\in\Delta(\mathcal{A}^{\otimes n-1})$ to denote the actions and the mixed strategies of the other players. We denote $U_{i}(x_{i},x_{-i}):=\mathbb{E}_{a_{i}\sim x_{i},a_{-i}\sim x_{-i}}\left[U_{i}(a_{i},a_{-i})\right].$
Learning protocol
We assume that the learner knows the game rule, and thus her own payoff function $U_{1}$. At every round $t$, all players take action simultaneously, and the learner only observes the opponents’ noisy actions $(a_{2}^{t},a_{3}^{t},\ldots,a_{n}^{t})$ that are sampled from their strategies.
@@
Nash Equilibrium (NE) is the most commonly-used solution concept for games: a mixed strategy $x\in\Delta(\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n})$ of all players is said to be NE if $x$ is a product distribution  [^4] The randomness in different players’ strategies are independent., and no player could gain by deviating from her own strategy while holding all other players’ strategies fixed. That is, for all $i\in[n]$ and $a^{\prime}_{i}\in\mathcal{A}_{i}$, $\mathbb{E}_{a\sim x}[U_{i}(a_{i},a_{-i})]\geq\mathbb{E}_{a\sim x}[U_{i}(a^{\prime}_{i},a_{-i})]$.
@@
There are also two equilibrium notions relaxing the notion of NE by no longer requiring $x$ to be a product distribution. It allows general joint distribution $x$ which describes correlated strategies among players. In particular, (1) $x$ is a Correlated Equilibrium (CE) if for all $i\in[n]$ and $a^{\prime}_{i}\in\mathcal{A}_{i}$, $\mathbb{E}_{a\sim x}[U_{i}(a_{i},a_{-i})\mid a_{i}]\geq\mathbb{E}_{a\sim x}[U_{i}(a^{\prime}_{i},a_{-i})\mid a_{i}]$, and (2) $x$ is Coarse Correlated Equilibrium (CCE) if for all $i\in[n]$ and $a^{\prime}_{i}\in\mathcal{A}_{i}$: $\mathbb{E}_{a\sim x}[U_{i}(a_{i},a_{-i})]\geq\mathbb{E}_{a\sim x}[U_{i}(a^{\prime}_{i},a_{-i})]$. The major difference between those two notions is in the cases when the agent deviates from her current strategy, whether she is still allowed to observe the randomness in drawing actions from the correlated strategy. The relationship among various equilibrium concepts is encapsulated by ${\rm NE}\subset{\rm CE}\subset{\rm CCE}$.
Two-player zero-sum games
@@
It is well-known that in two-player zero-sum games, all Nash equilibria share the unique payoff value 00. Furthermore, a Nash equilibrium is non-exploitable against any strategy that is not necessarily an equilibrium. In math, if $(\mu^{\star},\nu^{\star})$ is the Nash equilibrium, we have $\min_{\nu}U_{1}(\mu^{\star},\nu)=\max_{\mu}\min_{\nu}U_{1}(\mu,\nu)=0$.
Multiplayer or general-sum games
When the number of players is greater than two, or the games are no longer constant-sum games, Nash equilibria become PPAD-hard to compute  (Daskalakis et al., 2009) while CEs and CCEs can be still computed in polynomial time. All of these three concepts admit multiple equilibria with distinct payoffs. Furthermore, they no longer own strong guarantees, such as non-exploitness, when competing against non-equilibrium players.
2.2 Symmetric games and equal share
Symmetric games
For an $n$-player normal-form game with an action space $\{\mathcal{A}_{i}\}_{i=1}^{n}$ and a payoff $\{U_{i}\}_{i=1}^{n}$, we say the game is symmetric if (1) $\mathcal{A}_{i}=\mathcal{A}$, for all $i\in[n]$; (2) for any permutation $\sigma:[n]\rightarrow[n]$, we have $U_{i}(a_{1},\cdots,a_{n})=U_{\sigma^{-1}(i)}(a_{\sigma(1)},\cdots,a_{\sigma(n)})$.
In short, the payoffs of a symmetric game for employing a specific action are determined solely by the actions used by others, agnostic of the identities of the players using them. Thus, the payoff function of the first player denoted as $U_{1}$, is sufficient to encapsulate the entire game.
Symmetric games are popular in practice as they bring fairness among players. Technically, all asymmetric can be converted to symmetric games by randomizing the roles of the players at the beginning of the game. Nevertheless, casting the games in the symmetric form gives a natural and minimal baseline — the learning agent should attain an equal share in the long run.
Equal share
We say an agent attains an equal share, if the agent’s average payoff of the games is at least $C/n$ for a $n$-player symmetric constant-sum game with a total payoff of $C$.
It is not hard to see that shifting the total payoff of a game by any absolute constant will not alter the strategic aspects of the game. Therefore, without loss of generality, this paper sets a total payoff of $C=0$ and focuses on achieving an equal share in multiplayer symmetric zero-sum games.
3 Insufficiency of Equilibria and Self-play for Equal Share
In this section, we demonstrate that existing equilibria notions and the self-play from scratch algorithm are not sufficient to secure an equal share in multiplayer symmetric zero-sum games even under very basic settings. To illustrate this, we consider the following $3$-player majority vote game: